Calculate the kilojoules needed to heat 221 g of gold from 18 ∘C to 195 ∘C.
Answer #1

See the very important piece of information you need here is the specific heat of gold, since you have not provided it I am taking the standard specific heat of gold from Book, i.e. Cp = 0.129 J/goC

now coming to the main equation you need to get the amount of heat required is: Q = m * Cp * ΔT

where Q = required heat

m = mass of gold, ΔT = temperature increase

now mass of gold is given m = 221g, ΔT = (195 - 18) = 177 oC

hence Q = 221g * 0.129J/goC * 177oC

Q = 5046.093 J = 5.046 kJ

so the amount of heat required is 5.046 kilojoules